OpenStudy (anonymous):

If a=2b, 3b=4c, and 5c=6d, what is the ratio of a to d?

OpenStudy (anonymous):

a = 2b this means that b = a/2 3b = 4c this means 3a/2 = 4c or c = 3a/8 and so on.